OpenStudy (mallorysipp234):

A population of flies grows according to the function p(x) = 2(4)x, where x is measured in weeks. A local spider has set up shop and consumes flies according to the function s(x) = 2x + 5. What is the population of flies after two weeks with the introduced spider? 15 flies 23 flies 32 flies 36 flies

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (mallorysipp234):

so it's p(2) = 2(4)(2) right since x = the 2 weeks that we're trying to find right?

OpenStudy (mallorysipp234):

but 2 x 2 x 4 = 16. and that's none of the choices. so im confused.

OpenStudy (skullpatrol):

Are you sure this function: p(x) = 2(4)x is copied down correctly?

OpenStudy (mallorysipp234):

it's p(x) = 2(4)^x

OpenStudy (skullpatrol):

:)

OpenStudy (mallorysipp234):

32! oh okay i see it now. its 32 right???

OpenStudy (mallorysipp234):

can you help me with this one @skullpatrol @sourwing @donny471

OpenStudy (skullpatrol):